Hospital A

Hospital B

Hospital C

Patient days TYD:

20,000

20,000

20,000

Gross Dept expenses:

$700,000

$700,000

$700,000

YTD

Food Cost

YTD Sales & Charges

YTD

Food Cost

YTD Sales & Charges

YTD

Food Cost

YTD Sales & Charges

Floor stock, nourishments

$15,000

$0.00

$15,000

$15,000

$15,000

$15,000

Internal catering

$30,000

$0.00

$30,000

$30,000

$30,000

$75,000

Catering cash

$10,000

$ 25,000

$10,000

$ 25,000

$10,000

$ 25,000

Cafeteria cash

$68,000

$150,000

$68,000

$150,000

$68,000

$150,000

Vending cash

$10,000

$ 20,000

$10,000

$ 20,000

$10,000

$ 20,000

Note: YTD = Year-to-Date

2. Which hospital has the lowest revenue per patient day?
Answer #1

Ans:- In this question, we need to find the hospital that has the lowest revenue per patient day.

Patient day: - It is basically a time unit that is used to determine the cost to keep a patient for one day in the hospital.

First, we need to find the Total Revenue amount spend by Hospital A, B, and C.

Total Revenue Amount spend by the hospitals will be given by (Gross Dept Expenses + Total YTD Food cost + Total sales& charges

Total Revenue Spend by Hospital A

=$700000 + ( $15000 + $30000 +$10000 +$68000 + $10000) + ($25000 + $150000 + $20000) =$1028000

Revenue Spend by Hospital A per patient day will be given by

( Total Revenue Spend by A / Total unit of Patient day by hospital A)

= $1028000/20000 = $51.40

Revenue Spend by Hospital B

=$700000 +( $15000 + $30000 +$10000 +$68000 + $10000) + ( $15000 + $30000 + $25000 + $150000 + $20000) = $1073000.

Revenue Spend by Hospital B per patient day will be ( $1073000/20000) =$53.65

Revenue Spend by Hospital C

= $700000 +( $15000 + $30000 +$10000 +$68000 + $10000) +( $15000 +$75000+ $25000 + $150000 + $20000) =$1118000

Revenue Spent by Hospital C per patient day will be ($488000/20000) =$55.90

Therefore Hospital A has the lowest revenue on per patient day i.e $51.40